# Quantum Key Distribution via Charge Teleportation

### v1

@@ F.4 Bit-Flip Error
We now consider the effect of a local bit-flip error channel on the shared resource state. This error is modeled by the transformation $\rho\rightarrow(1-p)\rho+pX_{n}\rho X_{n}$, which applies a Pauli-$X$ gate to the qubit at site $n$ with probability $p$. Due to the local nature of the protocol’s key operators, it is sufficient to analyze the impact of such errors occurring at Alice’s site ($n=0$) and Bob’s site ($n=N$) [QKDbyQET].
@@ F.6 Bob site Phase-Flip Error
  • All plots show a strong, nearly linear degradation of the teleported signal as a function of the error probability $p$. This rapid decay is characteristic of an error that directly inverts the protocol’s intended outcome.

### v2

@@ F.4 Bit-Flip Error
We now consider the effect of a local bit-flip error channel on the shared resource state. This error is modeled by the transformation $\rho\rightarrow(1-p)\rho+pX_{n}\rho X_{n}$, which applies a Pauli-$X$ gate to the qubit at site $n$ with probability pp. Due to the local nature of the protocol’s key operators, it is sufficient to analyze the impact of such errors occurring at Alice’s site ($n=0$) and Bob’s site ($n=N$) [QKDbyQET].
@@ F.6 Bob site Phase-Flip Error
  • All plots show a strong, nearly linear degradation of the teleported signal as a function of the error probability pp. This rapid decay is characteristic of an error that directly inverts the protocol’s intended outcome.